OpenStudy (anonymous):

4z-10=3(6-8z) showing all work

OpenStudy (anonymous):

4z - 10 = 3(6 - 8z) 4z - 10 = 18 -24z 4z + 24z = 18 + 10 28z = 28 z = 28/28 z = 1